OpenStudy (anonymous):

MEDAL Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

my idea ?

OpenStudy (anonymous):

i really dont know i suck at math

OpenStudy (anonymous):

just the answer

OpenStudy (anonymous):

thank you, can you help me in 2 more

OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (charlotte123):

Sorry for coming In late! For the first question - The correct choice would be the last one ---> http://www.wolframalpha.com/input/?i=-2x+%2B+y+%3D+1+and+-4x+%2B+y+%3D+-1 <---- The second one Is rather confusing though >.>

OpenStudy (charlotte123):

@realmadridCR7 Hope that helped!